OpenStudy (anonymous):

Lisa typed a 1000-word essay at an average rate of 20 words per minute. If she started typing at 6:20 p.m. and did not take any breaks, at what time did Lisa finish typing the essay?

OpenStudy (anonymous):

= 1000/ 20 = 50 min => 6:20 + 50 = 7:10 pm

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

any questions?

OpenStudy (anonymous):

no thank you